OpenStudy (anonymous):

How many solutions exist to this system? 24x-9y=3 8x-3y=-3 A. 1 B. 2 C. None D. Infinite

OpenStudy (abb0t):

One

OpenStudy (anonymous):

24x - 9y = 3 8x - 3y = -3--->(-3)8x - 3y = -3 ------------------- 24x - 9y = 3 -24x +9y = 9 (result of multiplying by -3) ------------ 0 + 0 = 12 C. The variables cancel each other out so there is no solution.

OpenStudy (anonymous):

Thanks :-)

OpenStudy (anonymous):

welcome...thats what I am here for :)